Problem 9-03A Oriole Company had the following assets on January 1, 2022. Item Cost Purchase Date Useful Life (in years) Salvage Value Machinery $62,000 Jan. 1, 2012 10 $ 0 Forklift 21,000 Jan. 1, 2019 5 0 Truck 27,400 Jan. 1, 2017 8 3,000 During 2022, each of the assets was removed from service. The machinery was retired on January 1. The forklift was sold on June 30 for $11,100. The truck was discarded on December 31. Journalize all entries required on the above dates, including entries to update depreciation, where applicable, on disposed assets. The company uses straight-line depreciation. All depreciation was up to date as of December 31, 2021. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Debit Credit choose a transaction date enter an account title enter a debit amount enter a credit amount enter an account title enter a debit amount enter a credit amount choose a transaction date enter an account title to record depreciation expense on forklift enter a debit amount enter a credit amount enter an account title to record depreciation expense on forklift enter a debit amount enter a credit amount (To record depreciation expense on forklift) choose a transaction date enter an account title to record sale of forklift on June 30 enter a debit amount enter a credit amount enter an account title to record sale of forklift on June 30 enter a debit amount enter a credit amount enter an account title to record sale of forklift on June 30 enter a debit amount enter a credit amount enter an account title to record sale of forklift on June 30 enter a debit amount enter a credit amount (To record sale of forklift) choose a transaction date enter an account title to record depreciation expense on truck enter a debit amount enter a credit amount enter an account title to record depreciation expense on truck enter a debit amount enter a credit amount (To record depreciation expense on truck) choose a transaction date enter an account title to record sale of truck on December 31 enter a debit amount enter a credit amount enter an account title to record sale of truck on December 31 enter a debit amount enter a credit amount enter an account title to record sale of truck on December 31 enter a debit amount enter a credit amount (To record sale of truck) Click if you would like to Show Work for this question: Open Show Work
Answer #1

Solution

Oriole Company

Entries to record retirement of assets:

  1. Machinery –

Date

Account Titles

Debit

Credit

Jan 1, 2022

Accumulated Depreciation – Machinery

$62,000

Machinery

$62,000

(To record machinery retired)

Computations:

Cost = $62,000

Dt of purchase – jan 1, 2012

Useful life = 10 years

Salvage value =0

Machinery retired

Book value on date of retirement, Jan 1, 2022

St. line depreciation

Annual depreciation = 62,000/10 = 6,200

Since 10 years expired, the machinery is fully depreciated, leaving no book value.

Forklift –

Entry to record 6 months depreciation,

Date

Account Titles

Debit

Credit

June 30, 2022

Depreciation Expense

$2,100

Accumulated Depreciation

$2,100

Entry to record sale,

Date

Account Titles

Debit

Credit

June 30, 2022

Cash

$11,100

Accumulated Depreciation

$14,700

Forklift

$21,000

Gain on Sale

$4,800

(To record sale of forklift and resulting gain)

Computations:

Cost = $21,000

Dt of purchase – jan 1, 2019

Useful life = 5 years

Salvage value =0

Sold on June 30, 2022 for $11,100

Book value on date of sale, June 30, 2022

St. line depreciation

Annual depreciation = 21,000/5 = 4,200

Accumulated Depreciation till date of sale – 3 years depreciation + ½ year depreciation in 2022

= (3 x 4,200) + (4,200 x 6/12) = $14,700

Book value, June 30, 2022 = 21,000 – 14,700 = $6,300

Sale value = $11,100

Gain on sale = 11,100 – 6,300 = $4,800

  • Truck

Entry to record depreciation expense,

Date

Account Titles

Debit

Credit

Dec 31, 2022

Depreciation Expense

$3,050

Accumulated Depreciation

$3,050

Entry to record discard of truck,

Date

Account Titles

Debit

Credit

Dec 31, 2022

Accumulated Depreciation

$18,300

Loss on Disposal

$9,100

Truck

$27,400

(To record loss on disposal of truck)

Computations:

Cost = $27,400

Dt of purchase – jan 1, 2017

Useful life = 8 years

Salvage value =$3,000

Truck discarded on December 31

SL Depreciation = (27,400 – 3,000)/8 = $3,050

Years of use 6 years

Accumulated depreciation = 6 x 3,050 = $18,300

Book value on date of discard = 27,400 – 18,300 = $9,100        

Loss on discard = $9,100